OpenStudy (anonymous):

A man wishes to fence a rectangular lot. if the used material costing $1.50 a meter for the front of the lot and the material costing $1.30 a meter for the other 3 sides, the fence will cost $380. if he uses the cheaper material from all sides of the lot the fence will cost $364. Find the dimension of the lot.

hartnn (hartnn):

what equations u got ?

OpenStudy (anonymous):

x=width of the lot y=length of its lot (front and back of the lot) 1.30(2x)+1.30(y)+1.50(y)=380 1.30(2x)+1.30(2y)=364

hartnn (hartnn):

these are correct need help solving these?? i m getting 60,80.

hartnn (hartnn):

just subtract 2nd equation from 1st equation.

OpenStudy (anonymous):

i need a complete solution :)

OpenStudy (radar):

If you want the solition you have it in the third post above this post. Do you want the steps?

hartnn (hartnn):

but u are clear with equations right? i will give complete solution starting from the equations,ok??

OpenStudy (radar):

I think that might be helpful @hartnn

OpenStudy (anonymous):

ok.. i want a complete solution from the equation that i have given..:)

OpenStudy (anonymous):

thank you:)

hartnn (hartnn):

ok, 1.30(2x)+1.30(y)+1.50(y)=380---->(1) 1.30(2x)+1.30(2y)=364---------->(2) subtracting (2) from (1), 0.2y=16 y=80=length of lot substitute back in (2), 2.6x=364-2.6*80=156 x=156/2.6=60=width

hartnn (hartnn):

i skipped a step 2.8y-2.6y=380-364 0.2y=16

hartnn (hartnn):

ask if still have doubts somewhere.